How many moles of gas does it take to occupy 120 liters at a pressure of 2.3 atmospheres and a temperature of 340 K?
1
Expert's answer
2020-12-10T07:51:37-0500

pV=nRT

p = 2.3 atm = 202650 Pa

V = 120 l = 0.12 m3

T = 340 K

n = pV/(RT) = 202650*0.12/(8.314*340) = 8.60 mol